what volume will 2.5 miles of nitrogen occupy at a pressure of .725 atm and a temperature of 23°C
1
Expert's answer
2020-05-14T14:33:06-0400

Given,"P=0.725 \ atm,V=?,n=2.5,T=23\degree C=273+23=296K,R=0.0821"

Using ideal gas equation,"PV=nRT"

"0.725\times V=2.5\times 0.0821\times 296\implies V=\frac{2.5\times 0.0821\times 296}{0.725}" "=83.799\ L"